Suppose a firm's total revenues depends on the amount produced (q) according to the functoins TR=70q-q, total cost also depends on q; TC=q²+30q+100 a. What level of output should the firm produce in order to maximize profit? b. Show that the second order conditions for a maximum are satisfied at the out put level found in part (a). c. Does the solution calculated here obey the condition "marginal revenue equal marginal cost" rule? Explain.
1
Expert's answer
2021-10-26T17:23:15-0400

a.

"\pi=TR-TC\\=(70q-q^2)-(q^2+30q+!00)\\=70q-q^2-q^2-30q-100\\=40q-2q^2-100"

necessary condition will be as dollows

"\frac{\delta \pi}{\delta q}=0\\40-4q=0\\4q=40\\q=10"

therefore the profit will be

"\pi=TR-TC\\=40q-2q^2-100\\=40(10)-2(10)^2-100\\400-200-100\\=\$100"

10 units will be produced to get a profit of $100


b.

the second order is "\frac{\delta ^2\pi}{\delta q^2}=0"

"\frac{\delta ^2\pi}{\delta q^2}=0\\

\frac{\delta ^2(40-4q)}{\delta q^2}\leq0\\-4\leq0"


c.

"MR=\frac{d(TR)}{dq}\\=70-2q\\MC=\frac{d(TC)}{dq}\\=2q+30\\MR=MC\\70-2q+30\\4q=40\\q=10"

the solution obeys the rule of mrginal revenue equa;s to marginal cost